# Enhancing Oscillator-Based Ising Machine Models with Amplitude Dynamics and Polynomial Interactions

### v1

@@ IV.1 Model Evaluation
We compare our proposed Hamiltonian (Eq. (15)) against the previously established baseline [23] (Eq. (14)), which also incorporates the amplitude dynamics for oscillator-based Ising machines. We simulate these models by numerically solving their governing differential equations.
Overall Performance We evaluate model performance for cubic problems on Boolean Satisfiability Problems (SAT) from the SATLIB benchmark [71], using instances from the uf-20, uf-50, uf-75, uf-100 and uf-150 categories, totaling 1,700 instances. For each instance, 100 random initial conditions are generated.
Figure 1: Percentage of solvable instances across problem sets (UF-20 to UF-150) from SATLIB [71] under 100 runs per instance. The proposed Hopf model consistently outperforms the baseline model [23].
The results show that our approach consistently outperforms the baseline Hopf model, solving a higher percentage of instances and achieving lower final energy values more frequently. These improvements indicate superior optimization performance, as illustrated in Fig. 1 and Fig. 2.
Figure 2: Final energy distribution for solving 3-SAT problems across approximately 170,000 runs per model. The proposed Hopf model achieves lower final energies with higher frequency, demonstrating better performance in finding optimal or near-optimal solutions.
Energy minimization tendency Our proposed Hopf model has a clear real-valued Lyapunov function indicating the energy of the system. The derived oscillator dynamics also provides a clear energy minimization tendency.
Consider a simple example with three oscillators and the optimization target is $E=s_{1}s_{2}s_{3}$ where $s_{i}\in\{-1,+1\}$. Map the optimization problem to both of the models. As shown in Fig. 3, we observe distinct behaviors between the models. Neither the real part, imaginary part, nor the absolute value of the baseline Hopf oscillator model $\mathcal{H}$ decreases monotonically, while our proposed model will consistently seek the minimum of the energy.
Figure 3: Energy trajectories over time for the baseline model [23] and the proposed Hopf model.
@@ IV.2 Higher-order Models with Wirtinger Potentials
    Upon binarization, the energy function would optimally yield the same values as the discrete Ising formulation. For instance, with a complex oscillator representation $z$, one may map $\operatorname{Re}(z)>0$ to $s_{i}=+1$ and $\operatorname{Re}(z)<0$ to $s_{i}=-1$; alternatively, in a phase-only model, mapping $\phi_{i}=0$ to $s_{i}=+1$ and $\phi_{i}=\pi$ to $s_{i}=-1$ achieves this equivalence.
There are multiple ways to construct an energy function that satisfies these conditions. For simplicity, we focus on an oscillator system that involves only $m$ interactions of order $k$, assuming that all coupling coefficients are unity. In this case, the energy function can be expressed as follows:
$\mathcal{H}^{C}=\sum_{i=1}^{m}\prod_{j=1}^{k}\left(z^{i}_{j}+\left(z^{i}_{j}\right)^{*}\right)$ (22)
where $z^{i}_{j}$ denotes the complex representation of the $j$-th oscillator in the $i$-th interaction. We refer to this as the Complete Potential, or C-potential for short. However, a suitable subset of its terms can be selected to form a new energy function that still satisfies the desired conditions. When expanded, the products can be organized into $k+1$ groups of the form
$\mathcal{H}^{p}=\sum_{i}\left[\prod_{j=1}^{p}z^{i}_{j}\right]\left[\prod_{j=p+1}^{k}(z^{i}_{j})^{*}\right]$ (23)
where $p\in\{0,1,\dots,\lfloor{k/2}\rfloor\}$. Each $H^{p}$ is real-valued and has identical values for $\vec{z}\in\mathbb{R}^{n}$, preserving Property 1. We refer to the family of cost functions $\{\mathcal{H}^{p}\}$ as Wirtinger potentials for $\mathcal{H}$, with $\mathcal{H}^{p}$ referred to as the $p$-potential for $\mathcal{H}$.
For example, when $k=3$ and assuming unity coupling coefficients, the C-potential is
$\mathcal{H}^{C}=\sum_{i,j,k}\left(z_{i}+z_{i}^{*}\right)\left(z_{j}+z_{j}^{*}\right)\left(z_{k}+z_{k}^{*}\right).$ (24)
The 0-potential is then
$\mathcal{H}^{0}=\sum_{i,j,k}\left(z_{i}z_{j}z_{k}+z_{i}^{*}z_{j}^{*}z_{k}^{*}\right),$ (25)
which is manifestly real and preserves the Ising equivalence upon binarization. Alternatively, one may define the 1-potential as
$\mathcal{H}^{1}=\sum_{i,j,k}\left(z_{i}z_{j}z_{k}^{*}+z_{i}^{*}z_{j}^{*}z_{k}\right),$ (26)
which similarly fulfills the required conditions.
@@
Wirtinger potentials enable flexibility in modeling oscillator systems while maintaining consistency with the Ising paradigm. For example, in a 3-SAT problem, the optimization target comprises three types of terms: linear, quadratic, and cubic (corresponding to $k=1$, $2$, and $3$, respectively). This means that multiple oscillator modeling options exist for each type of term. The energy function Eq. (15) is one such option, and our empirical evaluations indicate that this combination delivers the best performance in our benchmarks. However, it may not be the optimal choice for all problem types. In the following section (IV.3), we discuss a real case that leverages the variety of these models to achieve effective binarization for oscillator-based Ising machines. A thorough analysis of higher-order systems would require extensive theoretical and experimental study, which is beyond the scope of this work. Future research may explore these formulations to further enhance the capabilities of oscillator-based Ising machines.
@@ IV.3 Binarization Behavior: Quadratic Problems
Recall that reading out solutions from oscillator-based Ising machines (OIMs) requires converting each oscillator’s continuous phase (from 00 to $2\pi$) into discrete Ising spins ($+1$ or $-1$). Phase conversion typically involves introducing a binarization axis $\vec{b}$, which separates phases into two groups. Denoting $z=x+iy$ in vector notation as $\vec{z}=(x,y)^{T}$, we define the binarization function $f_{B}(z)=\text{sign}(\vec{z}\times\vec{b})$. The most convenient choice for $\vec{b}$ is typically the imaginary axis, hence we have $f_{B}(x+iy)=\text{sign}(x)$. Fig. 4 illustrates three typical binarization behaviors. A system that has not binarized with respect to $\vec{b}$ will give an incorrect solution $f_{B}(\vec{z})$. The system has still bifurcated, however the axis of bifurcation $\vec{\mathfrak{b}}$ is not parallel to $\vec{b}$, causing erroneous spin assignments. Systems which are binarized with respect to $\vec{b}$ can differ in the magnitude of binarization. A weakly binarized system is susceptible to misclassifications owing to readout error, while a strongly binarized system is more robust to classification error. In this section, we explore the implications of our Wirtinger potential formalism for phase binarization.
Whether $\vec{\mathfrak{b}}=\vec{b}$ depends on the problem formulation. Consider a quadratic Hamiltonian $\mathcal{H}(s_{1},s_{2})=s_{1}s_{2}$. $\mathcal{H}$ has three main Wirtinger potentials, the 0-potential $\mathcal{H}^{0}(z_{1},z_{2})$ , the 1-potential $\mathcal{H}^{1}(z_{1},z_{2})$ , and their convex combination, the “complete” C-potential $\mathcal{H}^{C}(z_{1},z_{2})$. Their explicit forms are given by
$\begin{split}\mathcal{H}^{0}(z_{1},z_{2})&=z_{1}z_{2}+z^{*}_{1}z^{*}_{2}\\ \mathcal{H}^{1}(z_{1},z_{2})&=z_{1}^{*}z_{2}+z_{1}z^{*}_{2}\\ \mathcal{H}^{C}(z_{1},z_{2})&=0.5(z_{1}z_{2}+z^{*}_{1}z^{*}_{2})\\ &+0.5(z_{1}^{*}z_{2}+z_{1}z^{*}_{2}).\end{split}$ (27)
@@
$\begin{split}\mathcal{H}^{1}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=e^{i(\phi-\phi)}z_{1}^{*}z_{2}+e^{i(\phi-\phi)}z_{1}z^{*}_{2}\\ &=z_{1}^{*}z_{2}+z_{1}z^{*}_{2}=\mathcal{H}^{1}(z_{1},z_{2}).\end{split}$ (28)
@@
$\begin{split}\mathcal{H}^{0}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=e^{i2\phi}z_{1}z_{2}+e^{-i2\phi}z_{1}^{*}z_{2}^{*}\\ &\neq\mathcal{H}^{0}(z_{1},z_{2})\end{split}$ (29)
@@
$\begin{split}\mathcal{H}^{C}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=\gamma e^{i2\phi}z_{1}z_{2}+0.5e^{-i2\phi}z_{1}^{*}z_{2}^{*}\\ &+0.5(z_{1}^{*}z_{2}+z_{1}z_{2}^{*})\\ &\neq\mathcal{H}^{C}(z_{1},z_{2})\end{split}$ (30)
@@
$\begin{split}\mathcal{H}^{0}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=\frac{1}{2}[\cos(\pi+\pi/3)+\cos(\pi-\pi/3)]\\ &=-0.5\neq-1.0=\mathcal{H}^{0}(z_{1},z_{2}).\end{split}$ (31)
Investigating the Hessians provides more context. For convenience, we assume that the local confining potentials confine each oscillator to unit amplitude, allowing for a phase-reduced model
$\begin{split}\mathcal{H}^{0}(\theta_{1},\theta_{2})&=\cos(\theta_{1}+\theta_{2})\\ \mathcal{H}^{1}(\theta_{1},\theta_{2})&=\cos(\theta_{1}-\theta_{2})\\ \mathcal{H}^{C}(\theta_{1},\theta_{2})&=0.5\cos(\theta_{1}-\theta_{2})+(1-0.5)\cos(\theta_{1}+\theta_{2})\end{split}$ (32)
@@
$\begin{split}\nabla^{2}\mathcal{H}^{0}(\theta_{1},\theta_{2})&=\begin{bmatrix}-\cos(\theta_{1}+\theta_{2})&-\cos(\theta_{1}+\theta_{2})\\ -\cos(\theta_{1}+\theta_{2})&-\cos(\theta_{1}+\theta_{2})\end{bmatrix}\\ \nabla^{2}\mathcal{H}^{1}(\theta_{1},\theta_{2})&=\begin{bmatrix}-\cos(\theta_{1}-\theta_{2})&\cos(\theta_{1}-\theta_{2})\\ \cos(\theta_{1}-\theta_{2})&-\cos(\theta_{1}-\theta_{2})\end{bmatrix}\\ \nabla^{2}\mathcal{H}^{C}(\theta_{1},\theta_{2})&=0.5\nabla^{2}\mathcal{H}^{0}(\theta_{1},\theta_{2})\\ &+0.5\nabla^{2}\mathcal{H}^{1}(\theta_{1},\theta_{2})\end{split}$ (33)
@@
$\nabla^{2}\mathcal{H}^{1}(\theta_{1},\theta_{2})$ and $\nabla^{2}\mathcal{H}^{0}(\theta_{1},\theta_{2})$ have eigenvalues $\{0,2\}$, however their null spaces $\mathcal{N}$ are disjoint. Specifically, $\mathcal{N}(\nabla^{2}\mathcal{H}^{0})=\text{Span}\{(1,-1)^{T}\}$, $\mathcal{N}(\nabla^{2}\mathcal{H}^{1})=\text{Span}\{(1,1)^{T}\}$. In other words, the 1-potential has zero curvature along the $(1,1)^{T}$ vector while being strongly convex along $(1,-1)$. In contrast, the 0-potential is strongly convex along $(1,1)^{T}$ while having zero curvature along $(1,-1)$. Neither potential is strongly convex in the neighborhood of their fixed points, causing each to have an uncountable number of solutions.
However, the convex combination $\nabla^{2}\mathcal{H}^{C}(\theta_{1},\theta_{2})$ is full-rank and positive definite in the neighborhood of all $\{0,\pi\}$ minima, meaning it guarantees descent to the unique phase minimizer within each neighborhood.
@@
$\nabla^{2}\mathcal{H}^{\text{SHIL}}=\begin{bmatrix}-\sin(2\theta_{1})&0\\ 0&-\sin(2\theta_{2})\end{bmatrix}$ (34)
which makes $\nabla^{2}\{\mathcal{H}^{1}+\mathcal{H}^{\text{SHIL}}\}$ full rank and positive definite in the neighborhood of $\{0,\pi\}$. Therefore, we have two options for binarizing quadratic Hamiltonians:
@@
    Add $\mathcal{H}^{\text{SHIL}}$ to the problem Hamiltonian
@@
    Use $\mathcal{H}^{C}(z_{1},z_{2})$
Option 1 is typical in oscillator-based IMs [28, 66], however, our analysis offers an alternative. $\mathcal{H}^{C}(z_{1},z_{2})$ or, more generally, a convex combination of the 1-potential $\mathcal{H}^{1}(z_{1},z_{2})$ and the 0-potential $\mathcal{H}^{0}(z_{1},z_{2})$, creates a strongly convex function in the neighborhood of $\{0,\pi\}$ fixed points. In the next section, we explore this possibility to induce binarization in 3-SAT problems while maintaining high-quality solutions.

### v2

@@ IV.1 Model Evaluation
We compare our proposed Hamiltonian (Eq. 15) against the previously established baseline [51] (Eq. 14), which also incorporates the amplitude dynamics for oscillator-based Ising machines. We simulate these models by numerically solving their governing differential equations.
Overall Performance: We evaluate model performance on 1,700 uniform random 3-SAT benchmarks from SATLIB [103]. Each set, ufN-M, denotes instances with N variables and M clauses; we use uf20-91, uf50-218, uf75-325, uf100-430, and uf150-645, all at a clause-to-variable ratio of approximately 4.3. This ratio lies at the empirical phase transition in satisfiability, where random 3-SAT instances shift from mostly satisfiable to mostly unsatisfiable and where difficulty peaks [104]. We define a solvable instance as one where at least one of 100 independent runs (with different initial conditions) reaches zero energy within a sufficient simulation time ($T=136$), allowing both models to achieve their best performance on the benchmark.
Figure 1: Percentage of solvable instances across problem sets (UF20-91 to UF150-645) from SATLIB [103] under 100 runs per instance. Instances are classified as solvable if at least one run reaches the ground state ($\mathcal{H}=0$). The error bars denote the 99% bootstrap confidence intervals (computed over 10,000 resamples). The proposed Hopf model consistently outperforms the baseline model [51].
The following results show that our approach consistently outperforms the baseline Hopf model: We solve a higher percentage of instances and achieve lower final energy values more frequently. These improvements indicate superior optimization performance, as illustrated in Fig. 1 and Fig. 2.
Figure 2: Final energy distribution for solving 3-SAT problems across approximately 170,000 runs per model. The proposed Hopf model achieves lower final energies with higher frequency, as shown in the histogram (top). The cumulative plot (bottom) further demonstrates the improved performance, showing a higher percentage of trials for the proposed model successfully reaching optimal or near-optimal solutions.
@@ IV.2 Higher-order Models with Wirtinger Potentials
    Upon binarization, the energy function would optimally yield the same values as the discrete Ising formulation. For instance, with a complex oscillator representation $z$, one may map $\operatorname{Re}(z_{i})>0$ to $s_{i}=+1$ and $\operatorname{Re}(z_{i})<0$ to $s_{i}=-1$; alternatively, in a phase-only model, upon binarization, mapping $\phi_{i}=0$ to $s_{i}=+1$ and $\phi_{i}=\pi$ to $s_{i}=-1$ achieves this equivalence.
There are multiple ways to construct an energy function that satisfies these conditions. The most natural approach is to replace spin $s_{i}$ with $\frac{z_{i}+z_{i}^{*}}{2|z_{i}|}$. The resulting cubic terms of an energy function (ignoring the constant factor) will be
$\mathcal{H}^{C}_{3}(z)=\sum_{i,j,k}\left(z_{i}+z_{i}^{*}\right)\left(z_{j}+z_{j}^{*}\right)\left(z_{k}+z_{k}^{*}\right).$ (24)
We refer to this as the Complete potential (or C-potential for short) for cubic interactions, as it involves a complete set of cross terms. However, a family of systematic subsets of its terms can each satisfy the desired conditions. Taking cubic terms as an example:
$\mathcal{H}^{0}_{3}(z)$ $=\sum_{i,j,k}\left(z_{i}z_{j}z_{k}+z_{i}^{*}z_{j}^{*}z_{k}^{*}\right),$ (25)
$\mathcal{H}^{1}_{3}(z)$ $=\sum_{i,j,k}\left(z_{i}z_{j}z_{k}^{*}+z_{i}^{*}z_{j}^{*}z_{k}\right),$
are all manifestly real and fulfill the required conditions. We number this family of cost functions $\{\mathcal{H}^{0}_{k},\mathcal{H}^{1}_{k},...,\mathcal{H}^{\lfloor k/2\rfloor}_{k}\}$ and refer to them collectively as Wirtinger potentials for $\mathcal{H}$, with $\mathcal{H}^{p}_{k}$ referred to as the $p$-potential of order $k$ for $\mathcal{H}$. Here $p$ refers to the number of conjugate terms in the leading product of each conjugate pair. For instance, each pair of $\mathcal{H}^{1}_{4}(z)$ has the form $z_{i}z_{j}z_{k}z_{\ell}^{*}+z_{i}^{*}z_{j}^{*}z_{k}^{*}z_{\ell}$, while $\mathcal{H}^{2}_{4}(z)$ conjugate pairs have the form $z_{i}z_{j}z_{k}^{*}z_{\ell}^{*}+z_{i}^{*}z_{j}^{*}z_{k}z_{\ell}$. A general definition for Wirtinger potentials of arbitrary order can be found in Appendix C. Each $\mathcal{H}^{p}_{k}$ is real-valued and has identical values for $z\in\mathbb{C}^{n}$, preserving Property 1. So too are properly scaled linear combinations of $\mathcal{H}^{p}_{k}$, including the special case $\mathcal{H}^{C}_{k}=\frac{1}{\lfloor k/2\rfloor}\sum_{p}\mathcal{H}^{p}_{k}$.
@@
Wirtinger potentials enable flexibility in modeling oscillator systems while maintaining consistency with the Ising paradigm. For example, in a 3-SAT problem, the optimization target comprises three types of terms: linear, quadratic, and cubic (corresponding to $k=1$, 22, and $3$, respectively). This means that multiple oscillator modeling options exist for each type of term. The energy function Eq. 15 is one such option, and our empirical evaluations indicate that this combination delivers the best performance in our benchmarks. However, it may not be the optimal choice for all problem types. In the following section (IV.3), we discuss a real case that leverages the variety of these models to achieve effective binarization for oscillator-based Ising machines. A thorough analysis of higher-order systems would require extensive theoretical and experimental study, which is beyond the scope of this work. Future research may explore these formulations to further enhance the capabilities of oscillator-based Ising machines.
@@ IV.3 Binarization Behavior: Quadratic Problems
Recall that reading out solutions to QUBO problems from OIMs requires converting each oscillator’s continuous phase (from 0 to $2\pi$) into discrete Ising spins (+1+1 or 1-1). Two (somewhat independent) factors can impact the final results: ① the state of the oscillators just prior to the readout and ② the actual readout process:
  1. 1.
    The state: At the end of annealing, depending on the governing dynamics, the oscillators can strongly binarize or not. Clearly, a strongly binarized system (Fig. 3-c or d) is more robust to classification error than a weakly binarized system (Fig. 3-a or b).
  2. 2.
[…]
Our Wirtinger potentials have implications for both factors as we explore next.
Figure 3: Example phase plots, where a system can either be strongly/weakly binarized (which corresponds the variance of the phase distribution) and correctly/incorrectly binarized (which corresponds to classification errors from axis misalignment).
Whether $\vec{\mathfrak{b}}=\vec{b}$ depends on the problem formulation. Consider a quadratic Hamiltonian $\mathcal{H}(s_{1},s_{2})=s_{1}s_{2}$. $\mathcal{H}$ has three main Wirtinger potentials, the 0-potential $\mathcal{H}^{0}_{2}(z_{1},z_{2})$ , the 1-potential $\mathcal{H}^{1}_{2}(z_{1},z_{2})$ , and their convex combinations, including the “complete” C-potential $\mathcal{H}^{C}_{2}(z_{1},z_{2})$. Their explicit forms are given by
$\begin{split}\mathcal{H}^{0}_{2}(z_{1},z_{2})&=z_{1}z_{2}+z^{*}_{1}z^{*}_{2}\\ \mathcal{H}^{1}_{2}(z_{1},z_{2})&=z_{1}^{*}z_{2}+z_{1}z^{*}_{2}\\ \mathcal{H}^{C}_{2}(z_{1},z_{2})&=0.5(z_{1}z_{2}+z^{*}_{1}z^{*}_{2})\\ &+0.5(z_{1}^{*}z_{2}+z_{1}z^{*}_{2}).\end{split}$ (26)
@@
$\begin{split}\mathcal{H}^{1}_{2}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=e^{i(\phi-\phi)}z_{1}^{*}z_{2}+e^{i(\phi-\phi)}z_{1}z^{*}_{2}\\ &=z_{1}^{*}z_{2}+z_{1}z^{*}_{2}=\mathcal{H}^{1}_{2}(z_{1},z_{2}).\end{split}$ (27)
@@
$\begin{split}\mathcal{H}^{0}_{2}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=e^{i2\phi}z_{1}z_{2}+e^{-i2\phi}z_{1}^{*}z_{2}^{*}\\ &\neq\mathcal{H}^{0}_{2}(z_{1},z_{2})\end{split}$ (28)
@@
$\begin{split}\mathcal{H}^{C}_{2}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=0.5e^{i2\phi}z_{1}z_{2}+0.5e^{-i2\phi}z_{1}^{*}z_{2}^{*}\\ &+0.5(z_{1}^{*}z_{2}+z_{1}z_{2}^{*})\\ &\neq\mathcal{H}^{C}_{2}(z_{1},z_{2})\end{split}$ (29)
@@
$\begin{split}\mathcal{H}^{0}_{2}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=\frac{1}{2}[\cos(\pi+\pi/3)+\cos(\pi-\pi/3)]\\ &=-0.5\neq-1.0=\mathcal{H}^{0}_{2}(z_{1},z_{2}).\end{split}$ (30)
Investigating the Hessians provides more context. For convenience, we assume that the local dynamics confine each oscillator to unit amplitude, allowing for a phase-reduced model
$\begin{split}\mathcal{H}^{0}_{2}(\theta_{1},\theta_{2})&=\cos(\theta_{1}+\theta_{2})\\ \mathcal{H}^{1}_{2}(\theta_{1},\theta_{2})&=\cos(\theta_{1}-\theta_{2})\\ \mathcal{H}^{C}_{2}(\theta_{1},\theta_{2})&=\frac{1}{2}\cos(\theta_{1}-\theta_{2})+\frac{1}{2}\cos(\theta_{1}+\theta_{2})\end{split}$ (31)
@@
$\begin{split}\nabla^{2}\mathcal{H}^{0}_{2}(\theta_{1},\theta_{2})&=\begin{bmatrix}-\cos(\theta_{1}+\theta_{2})&-\cos(\theta_{1}+\theta_{2})\\ -\cos(\theta_{1}+\theta_{2})&-\cos(\theta_{1}+\theta_{2})\end{bmatrix}\\ \nabla^{2}\mathcal{H}^{1}_{2}(\theta_{1},\theta_{2})&=\begin{bmatrix}-\cos(\theta_{1}-\theta_{2})&\cos(\theta_{1}-\theta_{2})\\ \cos(\theta_{1}-\theta_{2})&-\cos(\theta_{1}-\theta_{2})\end{bmatrix}\\ \nabla^{2}\mathcal{H}^{C}_{2}(\theta_{1},\theta_{2})&=\frac{1}{2}\nabla^{2}\mathcal{H}^{0}_{2}(\theta_{1},\theta_{2})+\frac{1}{2}\nabla^{2}\mathcal{H}^{1}_{2}(\theta_{1},\theta_{2})\end{split}$ (32)
@@
$\nabla^{2}\mathcal{H}^{1}_{2}(\theta_{1},\theta_{2})$ and $\nabla^{2}\mathcal{H}^{0}_{2}(\theta_{1},\theta_{2})$ have eigenvalues $\{0,2\}$, however their null spaces $\mathcal{N}$ are disjoint. Specifically, $\mathcal{N}(\nabla^{2}\mathcal{H}^{0}_{2})=\text{Span}\{(1,-1)^{\top}\}$, $\mathcal{N}(\nabla^{2}\mathcal{H}^{1}_{2})=\text{Span}\{(1,1)^{\top}\}$. In other words, the 1-potential has zero curvature along the $(1,1)^{\top}$ vector while being strongly convex along (1,1)(1,-1). In contrast, the 0-potential is strongly convex along $(1,1)^{\top}$ while having zero curvature along (1,1)(1,-1). Neither potential is strongly convex in the neighborhood of their fixed points, causing each to have an uncountable number of solutions.
However, the convex combination $\nabla^{2}\mathcal{H}^{C}_{2}(\theta_{1},\theta_{2})$ is full-rank and positive definite in the neighborhood of all $\{0,\pi\}$ minima, meaning it guarantees descent to the unique phase minimizer within each neighborhood.
@@
$\nabla^{2}\mathcal{H}^{\text{SHIL}}_{2}=2\begin{bmatrix}\cos(2\theta_{1})&0\\ 0&\cos(2\theta_{2})\end{bmatrix}$, (33)
which makes $\nabla^{2}\{\mathcal{H}^{1}_{2}+\mathcal{H}^{\text{SHIL}}_{2}\}$ full rank and positive definite in the neighborhood of $\{0,\pi\}$. Therefore, we have two options for binarizing quadratic Hamiltonians:
@@
    Add $\mathcal{H}^{\text{SHIL}}_{2}$ to the problem Hamiltonian.
@@
    Use $\mathcal{H}^{C}_{2}(z_{1},z_{2})$.
Option 1 is typical in OIM literature [56, 98]. However, our analysis offers an alternative. $\mathcal{H}^{C}_{2}(z_{1},z_{2})$ or, more generally, a convex combination of the 1-potential $\mathcal{H}^{1}_{2}(z_{1},z_{2})$ and the 0-potential $\mathcal{H}^{0}_{2}(z_{1},z_{2})$ creates a strongly convex function in the neighborhood of $\{0,\pi\}$ fixed points. In the next section, we explore this possibility to induce binarization in 3-SAT problems while maintaining high-quality solutions.